OpenStudy (anonymous):

What are the excluded values? y+5/y^2+4y-32

OpenStudy (anonymous):

its, . y = 4 and -8

OpenStudy (anonymous):

@cillawilla

OpenStudy (anonymous):

they need the excluded values right? It would be either, that what i said or -8 , 4

OpenStudy (anonymous):

y^2 + 4y - 32 = 0 (y + 8)(y - 4) = 0 y + 8 = 0 or y - 4 = 0 y = -8...........y = 4 The excluded values are -8 and 4.

OpenStudy (blank ):

Or it could look like this: http://www4c.wolframalpha.com/Calculate/MSP/MSP988220e529431652hga400001b172beeibc1fedb?MSPStoreType=image/gif&s=54&w=89.&h=40. You can't know for sure until the person asking the question clears it up a little.

OpenStudy (anonymous):

yeah but they are off line. so guess what

OpenStudy (anonymous):

We are both right! :D

OpenStudy (blank ):

I didn't give a solution yet, I only brought up a question.

OpenStudy (anonymous):

we dont know what they are asking until more in depth description.

OpenStudy (blank ):

Yep ^